Evaluate.
r + s/7 for r =51 and s = 12
r + s/7 =______. (Simplify)

Answer by richwmiller(17219)

we have the expression r+s/7
and we have values for r and s
so we substitute the values for the variables (r and s)
51+12/7
now we have to combine the improper fraction with the whole number 51.
we can do that several ways.
we can convert 12/7 to a whole number plus a fraction and then add it to 51
or we can convert 51 to an improper fraction also over 7 and add the two numerators and then reduce the improper fraction
51=51*7/7
(51*7+12)/7=(357+12)/7=369/7=52 5/7
the other way would have been convert 12/7 to 1 5/7 and then add 1 5/7 to 51 to get 52 5/7